What is the answer to 11/2x - 4/5x = 47/10
Answer by rfer(16322)

11/2x-4/5x=47/10
find common denominator
55/10x-8/10x=47/10
47/10x=47/10
x=(47/10)/(47/10)
x=47/10*10/47
x=470/470
x=1